OpenStudy (anonymous):

x+3/(x+1)(x+7) partial fraction..

OpenStudy (anonymous):

Apply the method for the other problem.

OpenStudy (anonymous):

\[ \frac{x+3}{(x+1) (x+7)}=\frac{A}{x+1}+\frac{B}{x+7} \]

OpenStudy (anonymous):

To find A, you multiply both sides by what?